# SMP: Reusable Score-Matching Motion Priors for Physics-Based Character Control

### v2

@@ Abstract.
Data-driven motion priors that can guide agents toward producing naturalistic behaviors play a pivotal role in creating life-like virtual characters. Adversarial imitation learning has been a highly effective method for learning motion priors from reference motion data. However, adversarial priors, with few exceptions, need to be retrained for each new controller, thereby limiting their reusability and necessitating the retention of the reference motion data when training on downstream tasks. In this work, we present Score-Matching Motion Priors (SMP), which leverages pre-trained motion diffusion models and score distillation sampling (SDS) to create reusable task-agnostic motion priors. SMPs can be pre-trained on a motion dataset, independent of any control policy or task. Once trained, SMPs can be kept frozen and reused as general-purpose reward functions to train policies to produce naturalistic behaviors for downstream tasks. We show that a general motion prior trained on large-scale datasets can be repurposed into a variety of style-specific priors. Furthermore SMP can compose different styles to synthesize new styles not present in the original dataset. Our method produces high-quality motion comparable to state-of-the-art adversarial imitation learning methods through reusable and modular motion priors. We demonstrate the effectiveness of SMP across a diverse suite of control tasks with physically simulated humanoid characters. Video demo available at https://youtu.be/ravlZJteS20
@@ 1. Introduction
Creating virtual characters that move with natural and life-like behaviors is fundamental to immersive digital experiences in animation, films, games, and virtual reality (VR) applications. While motion capture and procedural animation techniques can produce high-quality movements, the challenge lies in developing controller that enable physically simulated characters to exhibit similarly natural behaviors dynamically, in response to diverse tasks and environmental contexts. Traditional physics-based controllers trained without reference motion data often result in visually unnatural movements, lacking the subtle qualities of life-like motion (yin2007simbicon; hodgins1995animating; 2010-TOG-gbwc). Tracking-based methods improve realism by following reference motion clips, but they typically require the controller to rigidly mimic target motions frame-by-frame (peng2018deepmimic; liu2017learning; won2017train), limiting its flexibility to deviate from the reference and adapt behaviors to perform new tasks. Alternatively, distribution-matching methods, such as adversarial imitation learning (peng2021amp; ho2016generative), provide a more versatile approach for imitating motion data. Adversarial methods can learn flexible motion priors from motion dataset, which can act as task-agnostic measures of motion naturalness, allowing policies to produce behaviors that resemble natural motions across different tasks. However, adversarial methods typically require a prior (i.e. discriminator) to be trained jointly with a given policy, which limits the reusability and modularity of the learned priors. For each new policy, the prior must be continuously trained with data from the policy and data from the original reference dataset. As a result, the original dataset must be retained for perpetuity.
We contend that an ideal motion prior should be modular and reusable:
@@
  • Reusable: Once constructed, the same prior should be applicable across diverse tasks and multiple policies without retraining.
To achieve these criteria, we propose Score-Matching Motion Priors (SMP): a method for constructing reusable, modular motion priors that can be utilized to train control policies to produce naturalistic behaviors across diverse tasks. Given an unstructured motion dataset, we first train a motion diffusion model to capture the underlying data distribution independent of any task or control policy. Once trained, the diffusion model is kept frozen and repurposed as a prior via score distillation sampling (SDS) (poole2022dreamfusion), providing a robust similarity measure between simulated motions and the behaviors in the reference dataset. By reusing the learned SMP as a general style reward, our system enables a prior to be used to train new policies to perform a diverse suite of tasks with natural life-like behaviors.
The central contribution of this work is a method for constructing modular and reusable motion priors for physics-based character animation by combining reinforcement learning with score distillation. Our framework produces high-quality motions comparable to state-of-the-art adversarial imitation learning methods, while substantially improving modularity and reusability. SMP enables our system to completely discard the reference dataset during policy training. We show that a score-matching motion prior can be constructed from a task-agnostic diffusion model pretrained on large-scale motion datasets. This general-purpose motion prior can then be repurposed into style-specific priors through prompting and guidance. Furthermore, we show that priors for different styles can be composed to create new behavioral styles that are not present in the original dataset.
@@ 2. Related Work
Developing embodied agents that can act and react with life-like motions is essential for creating immersive experiences in games and virtual reality applications. This capability is also of vital importance in robotics, where naturalistic behaviors can improve safety, energy efficiency (escontrela2022adversarial), and the learning of broadly applicable skills from human data (Grauman_2022_CVPR). Given the expressive and nuanced nature of human motion, data-driven approaches have emerged as an effective paradigm for producing realistic, life-like behaviors by learning from reference motion data. Kinematics-based methods typically train models, such as autoregressive models (holden2016deep; zhang2018mode; holden2017phase), variational autoencoders (VAEs) (ling2020character; rempe2021humor; starke2024categorical), or diffusion models (zhang2022motiondiffuse; tevet2023human; shi2024interactive), on large motion datasets to synthesize plausible character animations. However, most of these methods do not explicitly enforce physical constraints, often leading to physically implausible behaviors, especially for new scenarios and tasks.
@@
In contrast, physics-based methods aim to create control policies that generate motion within simulated environments, governed by dynamic equations and realistic physical laws (wampler2014generalizing; raibert1991animation). These controllers are often constructed via trajectory optimization techniques or reinforcement learning (RL) (peng2017deeploco; wang2009optimizing). Previous data-free approaches based on heuristics, such as SIMBICON (yin2007simbicon), can achieve mechanically functional behaviors, but often produce unnatural motions. Manually designing heuristics that capture the expressiveness of real human movement remains a significant challenge (faloutsos2001composable; hodgins1995animating; witkin1988spacetime; 2010-TOG-gbwc). Instead of relying on hand-crafted heuristics, data-driven approaches based on model predictive control (MPC) or trajectory optimization have shown promise in emulating naturalistic human motions (sharon2005synthesis; muico2009contact; lee2010data; liu2010sampling). More recently, reinforcement learning-based motion tracking methods have enabled controllers to effectively imitate a wide range of reference motion clips, resulting in more life-like and agile behaviors (peng2018deepmimic; liu2017learning; won2017train). However, tracking-based methods limit a controller to closely following a given motion clip, which can impede the controller’s ability to generalize to new tasks, particularly when task objectives are not closely aligned with the reference motions. To address this limitation, many systems incorporate task-specific motion planners (bergamin2019drecon; liu2012terrain; park2019learning), or high-level policies (yao2022controlvae; yao2024moconvq; zhu2023neural), which select appropriate reference motions or latent-space skills for the controller to perform in order to complete a given task.
Distribution matching offers an alternative to motion tracking by training controllers to imitate the broader behavioral distribution of a motion dataset rather than tracking reference motions frame-by-frame. Generative Adversarial Imitation Learning (GAIL) approximates this objective using a learned discriminator to distinguish agent’s motions from dataset motions (ho2016generative). Adversarial Motion Priors (AMP) extend this idea to model flexible style objectives, enabling controllers to produce life-like behaviors for novel tasks that are not observed in the original dataset (peng2021amp). However, adversarial priors must be trained jointly with a specific policy for each new task, often requiring retraining and persistent access to the dataset. In contrast, our method achieves comparable performance to state-of-the-art adversarial imitation learning approaches by introducing a reusable and modular score-matching motion prior, which can even be shaped into stylistic priors beyond those contained in the original dataset.
@@
Diffusion models’ state-of-the-art performance across a wide range of domain has spurred growing interest in leveraging diffusion models for control. Given their ability to generate high-fidelity motion, a straightforward application is to use task-oriented diffusion models as motion planners (xu2025parc; tevet2024closd; serifi2024robot; ren2023insactor). These methods leverage auto-regressive motion diffusion models to predict target future trajectories, which are then executed by a low-level tracking controller. In addition to their use as motion planners, diffuse models have also been used to directly model controllers, enabling controllers to produce flexible multi-modal behaviors for tasks such as manipulation (chi2023diffusionpolicy; janner2022planning), and character control (huang2025diffuse; truong2024pdp; wu2025uniphys; Huang2024DiffuseLocoRL). Unlike these prior methods that focus on incorporating diffusion models as components of a controller, our work aims to repurpose pretrained diffusion models as task-agnostic behavioral priors within an optimization objective. Our method utilizes diffusion models as a reward function that can be integrated into general reinforcement learning frameworks to guide policy learning via score distillation sampling.
@@
Pretrained diffusion models are not only capable of generating samples via a denoising process, but can also serve as optimization objectives through score distillation sampling (SDS) (poole2022dreamfusion). SDS has enabled pretrained image and video diffusion models to be adapted for a wide range of downstream generative modeling tasks (wang2023prolificdreamer; Jiang2024Animate3D; Yin_2024_CVPR). Inspired by the success of SDS in the vision domain, recent works have also explored incorporating diffusion-based priors into reinforcement learning frameworks for control. A series of methods replace the discriminator in a GAIL-style setup with a diffusion model (Wang2024DiffAIL; NEURIPS2024DiffRewardAIL; Pang2025DPR; NEURIPS2024DIO), leveraging its expressive capabilities to differentiate between real and fake trajectories. However, these methods still rely on adversarial training and typically require continual updates to the diffusion model during policy optimization. Beyond adversarial frameworks, other systems have attempted to apply SDS-like techniques directly to policy learning. For example, luo2024textaware guide policy training using pretrained image or video diffusion models conditioned on text prompts. While these methods can produce behaviors that align with textual instructions, the resulting motions often appear unnatural. The work that is most reminiscent to ours is SMILING (wu2025diffusing), which train controllers using a score-matching objective similar to variational score distillation (VSD)  (wang2023prolificdreamer). However, unlike SMILING, which requires training task-specific diffusion models, our method trains a task-agnostic motion prior from unstructured motion data. This prior can then be reused to train diverse control policies by combining it with separate task objectives in a goal-conditioned reinforcement learning framework. We further introduce several key design decisions that enable high-quality naturalistic motions across a diverse repertoire of tasks.
@@ 3.1. Reinforcement Learning
Our physics-based controllers are trained through a reinforcement learning (RL) framework, where an agent interacts with an environment according to a policy $\pi$ in order to optimize a given objective $J(\pi)$ (sutton1998reinforcement). At each time step $t$, the agent observes the current state ${\mathbf{s}}_{t}$, then samples and executes an action according to the policy ${\mathbf{a}}_{t}\sim\pi({\mathbf{a}}_{t}|{\mathbf{s}}_{t})$. The environment then transitions to a new state according to the dynamics ${\mathbf{s}}_{t+1}\sim p({\mathbf{s}}_{t+1}|{\mathbf{s}}_{t},{\mathbf{a}}_{t})$, and the agent receives a scalar reward $r_{t}=r({\mathbf{s}}_{t},{\mathbf{a}}_{t},{\mathbf{s}}_{t+1})$. The objective is to learn a policy that maximizes the expected discounted return:
@@ 3.2. Diffusion Models and Score Distillation Sampling
In score-based generative modeling methods (song2019generative; song2021scorebased), a neural network $f:{\mathbb{R}}^{D}\rightarrow{\mathbb{R}}^{D}$ is trained to estimate the score of a point $\mathbf{x}$} under a given distribution $p({\mathbf{x}})$, which is defined as the gradient of the log-density $\nabla_{{\mathbf{x}}}\log p({\mathbf{x}})$. However, the predicted scores may be unreliable, as the available training data typically do not cover the entire space ${\mathbb{R}}^{D}$. In regions with low data density, score estimates tend to be inaccurate. To address this issue, vincent2011connection and song2021scorebased propose adding noise to the data. When the noise level is sufficiently large, the perturbed data distribution covers the entire space, enabling the training of more robust and scalable score estimators. This principle forms the foundation of our approach. Score-based generative modeling with multi-level noise perturbation allows the construction of reusable motion priors from relatively limited data that lies in a low-dimensional manifold.
@@
Diffusion models approximate a data distribution by progressively corrupting and subsequently denoising samples through a sequence of transformations (ho2020denoising). Given a clean sample from the data distribution, the forward process iteratively adds Gaussian noise over $N$ steps to form a Markov chain $\{{\mathbf{x}}^{i}\}_{i=0}^{N}$, defined as:
@@
without performing the full iterative diffusion process. Here, $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$ and $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. This forward diffusion process resembles the multi-level noise perturbation used in score-based generative models (song2021scorebased). To sample from the learned distribution $p({\mathbf{x}})$, a denoising network $f$ is typically trained using the simple DDPM objective (ho2020denoising):
@@
where ${\mathbf{x}}^{i}$ is obtained using Equation 3.
@@
Once a diffusion model $f$ is trained, samples can be generated by applying the reverse diffusion process or through score distillation sampling (SDS) (poole2022dreamfusion). SDS minimizes the KL divergence between the distribution of diffused samples derived from the forward diffusion process $q({\mathbf{x}}^{i}\mid{\mathbf{x}}^{0})$ and the reference data distribution learned by the pretrained diffusion model $f({\mathbf{x}}^{i})$. The gradient of the KL divergence can be estimated using the difference between the score from the forward diffusion process and the prediction from the diffusion model:
@@
where $\mathbf{x}$} is the sample being optimized, and $w(i)$ are coefficients determined by the diffusion noise schedule. Previous work has shown that the weighting function $w(i)$ has limited impact and can be substituted with uniform weighting without negatively affecting performance (threestudio2023). The SDS gradient can also be derived from the diffusion loss in Equation 4, omitting the Jacobian with respect to the score estimator $f$. The SDS loss can be simplified as
@@
Figure 2. Schematic overview of the system. The dashed arrows indicate components used only during policy training. A pretrained motion diffusion model serves as a reusable reward model for motion naturalness via score distillation sampling. The model can be style-conditioned, enabling the policy to learn specific skills or styles without retraining or continuous access to original motion data.

### v3

@@ Abstract.
Data-driven motion priors that can guide agents toward producing naturalistic behaviors play a pivotal role in creating life-like virtual characters. Adversarial imitation learning has been a highly effective method for learning motion priors from reference motion data. However, adversarial priors, with few exceptions, need to be retrained for each new controller, thereby limiting their reusability and necessitating the retention of the reference motion data when applied to downstream tasks. In this work, we present Score-Matching Motion Priors (SMP), which leverages pre-trained motion diffusion models and score distillation sampling (SDS) to create reusable task-agnostic motion priors. SMPs can be pre-trained on a motion dataset, independent of any control policy or task. Once trained, SMPs can be kept frozen and reused as general-purpose reward functions to train new policies to produce naturalistic behaviors for downstream tasks. We show that a general motion prior trained on large-scale datasets can be repurposed into a variety of style-specific priors. Furthermore, SMP can compose different styles to synthesize new styles not present in the original dataset. Our method can create reusable and modular motion priors that produce high-quality motions comparable to state-of-the-art adversarial imitation learning methods. In our experiments, we demonstrate the effectiveness of SMP across a diverse suite of control tasks with physically simulated humanoid characters. Video available at youtu.be/jBA2tWk6vzU.
@@ 1. Introduction
Creating virtual characters that move with natural and life-like behaviors is fundamental to immersive digital experiences in animation, films, games, and virtual reality (VR) applications. While motion capture and procedural animation techniques can produce high-quality movements, the challenge lies in developing controller that enable physically simulated characters to exhibit similarly natural behaviors dynamically, in response to diverse tasks and environmental contexts. Traditional physics-based controllers trained without reference motion data often result in visually unnatural movements, lacking the subtle qualities of life-like motions (Yin et al., 2007; Hodgins et al., 1995; Coros et al., 2010). Tracking-based methods improve realism by following reference motion clips, but they typically require controllers to rigidly mimic target motions frame-by-frame (Peng et al., 2018a; Liu and Hodgins, 2017; Won et al., 2017), limiting their flexibility to deviate from the reference and adapt behaviors to perform new tasks. Alternatively, distribution-matching methods, such as adversarial imitation learning (Peng et al., 2021; Ho and Ermon, 2016), provide a more versatile approach for imitating motion data. Adversarial methods can learn flexible motion priors from motion dataset, which can act as task-agnostic measures of motion naturalness, allowing policies to produce behaviors that resemble natural motions across different tasks. However, adversarial methods typically require a prior (i.e. discriminator) to be trained jointly with a given policy, which limits the reusability and modularity of the learned priors. For each new policy, the prior must be continuously trained with data from the policy and data from the original reference dataset. As a result, the original dataset must be retained for perpetuity.
We contend that for many practical control applications, an ideal motion prior should be modular and reusable:
@@
  • Reusable: Once constructed, the same prior should be applicable across diverse tasks and multiple policies without requiring further training.
To develop motion priors that satisfy these criteria, we propose Score-Matching Motion Priors (SMP): a method for constructing reusable, modular motion priors that can be utilized to train control policies to produce naturalistic behaviors across diverse tasks. Given an unstructured motion dataset, we first train a motion diffusion model to capture the underlying data distribution independent of any task or control policy. Once trained, the diffusion model is kept frozen and repurposed as a prior via score distillation sampling (SDS) (Poole et al., 2022), providing a robust similarity measure between simulated motions and the behaviors in the reference dataset. By reusing the learned SMP as a general style reward, our system enables a prior to be used to train new policies to perform a diverse suite of tasks with natural life-like behaviors.
The central contribution of this work is a method for constructing modular and reusable motion priors for physics-based character animation by leveraging diffusion models as reward models in reinforcement learning. Our framework produces high-quality motions comparable to state-of-the-art adversarial imitation learning methods, while substantially improving modularity and reusability. SMP enables our system to completely discard the reference dataset during policy training. We show that a score-matching motion prior can be constructed from a task-agnostic diffusion model pretrained on large-scale motion datasets. This general-purpose motion prior can then be repurposed into style-specific priors through prompting and guidance. Furthermore, we show that priors for different styles can be composed to create new behavioral styles that are not present in the original dataset. Code is available in MimicKit.
@@ 2. Related Work
Developing embodied agents that can act and react with life-like motions is essential for creating immersive experiences in games and virtual reality applications. This capability is also of vital importance in robotics, where naturalistic behaviors can improve safety, energy efficiency (Escontrela et al., 2022), and the learning of general-purposes skills from human data (Grauman et al., 2022). Given the expressive and nuanced nature of human motion, data-driven approaches have emerged as an effective paradigm for producing realistic, life-like behaviors by learning from reference motion data. Kinematics-based methods typically train models, such as autoregressive models (Holden et al., 2016; Zhang et al., 2018; Holden et al., 2017), variational autoencoders (VAEs) (Ling et al., 2020; Rempe et al., 2021; Starke et al., 2024), or diffusion models (Zhang et al., 2022; Tevet et al., 2023; Shi et al., 2024), on large motion datasets to synthesize plausible character animations. However, most of these methods do not explicitly enforce physical constraints, often leading to physically implausible behaviors, especially for new scenarios and tasks.
@@
In contrast, physics-based methods aim to create control policies that generate motion within simulated environments, governed by dynamic equations and realistic physical laws (Wampler et al., 2014; Raibert and Hodgins, 1991). These controllers are often constructed via trajectory optimization techniques or reinforcement learning (RL) (Peng et al., 2017; Wang et al., 2009). Previous data-free approaches based on heuristics, such as SIMBICON (Yin et al., 2007), can achieve mechanically functional behaviors, but often produce unnatural motions. Manually designing heuristics that capture the expressiveness of real human movement remains a significant challenge (Faloutsos et al., 2001; Hodgins et al., 1995; Witkin and Kass, 1988; Coros et al., 2010). Instead of relying on hand-crafted heuristics, data-driven approaches based on model predictive control (MPC) or trajectory optimization have shown promise in emulating naturalistic human motions (Sharon and van de Panne, 2005; Muico et al., 2009; Lee et al., 2010; Liu et al., 2010). More recently, reinforcement learning-based motion tracking methods have enabled controllers to effectively imitate a wide range of reference motion clips, resulting in more life-like and agile behaviors (Peng et al., 2018a; Liu and Hodgins, 2017; Won et al., 2017). However, tracking-based methods limit a controller to closely following a given motion clip, which can impede the controller’s ability to generalize to new tasks, particularly when task objectives are not closely aligned with the reference motions. To address this limitation, many systems incorporate task-specific motion planners (Bergamin et al., 2019; Liu et al., 2012; Park et al., 2019), or high-level policies (Yao et al., 2022, 2024; Zhu et al., 2023), which select appropriate reference motions or latent-space skills for the controller to perform in order to complete a given task.
Distribution matching offers an alternative to motion tracking by training controllers to imitate the broader behavioral distribution of a motion dataset rather than tracking reference motions frame-by-frame. Generative Adversarial Imitation Learning (GAIL) approximates this objective using a learned discriminator to distinguish agent’s motions from dataset motions (Ho and Ermon, 2016). Adversarial Motion Priors (AMP) extend this idea to model flexible style objectives, enabling controllers to produce life-like behaviors for novel tasks that are not observed in the original dataset (Peng et al., 2021). However, adversarial priors must be trained jointly with a specific policy for each new task, often requiring retraining and persistent access to the dataset. In contrast, our method achieves comparable performance to state-of-the-art adversarial imitation learning approaches by introducing a reusable and modular score-matching motion prior, which can even be shaped into stylistic priors beyond those contained in the original dataset.
@@
Diffusion models’ state-of-the-art performance across a wide range of domain has spurred growing interest in leveraging diffusion models for control. Given their ability to generate high-fidelity motion, a straightforward application is to use task-oriented diffusion models as motion planners (Xu et al., 2025; Tevet et al., 2024; Serifi et al., 2024; Ren et al., 2023). These methods leverage auto-regressive motion diffusion models to predict target future trajectories, which are then executed by a low-level tracking controller. In addition to their use as motion planners, diffuse models have also been used to directly model controllers, enabling controllers to produce flexible multi-modal behaviors for tasks such as manipulation (Chi et al., 2023; Janner et al., 2022), and character control (Huang et al., 2025; Truong et al., 2024; Wu et al., 2025b; Huang et al., 2024b). Unlike these prior methods that focus on incorporating diffusion models as components of a controller, our work aims to repurpose pretrained diffusion models as task-agnostic behavioral priors within an optimization objective. Our method utilizes diffusion models as a reward function that can be integrated into general reinforcement learning frameworks to guide policy learning via score distillation sampling.
@@
Pretrained diffusion models are not only capable of generating samples via a denoising process, but can also serve as optimization objectives through score distillation sampling (SDS) (Poole et al., 2022). SDS has enabled pretrained image and video diffusion models to be adapted for a wide range of downstream generative modeling tasks (Wang et al., 2023; Jiang et al., 2024; Yin et al., 2024). Inspired by the success of SDS in the vision domain, recent works have also explored incorporating diffusion-based priors into reinforcement learning frameworks for control. A series of methods replace the discriminator in a GAIL-style setup with a diffusion model (Wang et al., 2024; Lai et al., 2024; Pang et al., 2025; Huang et al., 2024a), leveraging its expressive capabilities to differentiate between real and fake trajectories. However, these methods still rely on adversarial training and typically require continual updates to the diffusion model during policy optimization. Beyond adversarial frameworks, other systems have attempted to apply SDS-like techniques directly to policy learning. For example, Luo et al. (2024) guide policy training using pretrained image or video diffusion models conditioned on text prompts. While these methods can produce behaviors that align with textual instructions, the resulting motions often appear unnatural. The work that is most reminiscent to ours is SMILING (Wu et al., 2025a), which train controllers using a score-matching objective similar to variational score distillation (VSD)  (Wang et al., 2023). However, unlike SMILING, which requires training task-specific diffusion models, our method trains a task-agnostic motion prior from unstructured motion data. This prior can then be reused to train diverse control policies by combining it with separate task objectives in a goal-conditioned reinforcement learning framework. We further introduce several key design decisions that enable high-quality naturalistic motions across a diverse repertoire of tasks.
@@ 3. Background
Our system leverages reinforcement learning to train physics-based control policies and diffusion models to construct score-matching motion priors used as training objectives. In this section, we review the core concepts of reinforcement learning and diffusion models that underpin our method.
@@ 3.1. Reinforcement Learning
Our physics-based controllers are trained through a reinforcement learning (RL) framework, where an agent interacts with an environment according to a policy $\pi$ in order to optimize a given objective $J(\pi)$ (Sutton et al., 1998). At each time step $t$, the agent observes the current state ${\mathbf{s}}_{t}$, then samples and executes an action according to the policy ${\mathbf{a}}_{t}\sim\pi({\mathbf{a}}_{t}|{\mathbf{s}}_{t})$. The environment then transitions to a new state according to the dynamics ${\mathbf{s}}_{t+1}\sim p({\mathbf{s}}_{t+1}|{\mathbf{s}}_{t},{\mathbf{a}}_{t})$, and the agent receives a scalar reward $r_{t}=r({\mathbf{s}}_{t},{\mathbf{a}}_{t},{\mathbf{s}}_{t+1})$. The objective is to learn a policy that maximizes the expected discounted return:
@@ 3.2. Diffusion Models and Score Distillation Sampling
In score-based generative modeling methods (Song and Ermon, 2019; Song et al., 2021), a neural network $f:{\mathbb{R}}^{D}\rightarrow{\mathbb{R}}^{D}$ is trained to estimate the score of a point $\mathbf{x}$} under a given distribution $p({\mathbf{x}})$, which is defined as the gradient of the log-density $\nabla_{{\mathbf{x}}}\log p({\mathbf{x}})$. However, the predicted scores may be unreliable, as the available training data typically do not cover the entire space ${\mathbb{R}}^{D}$. In regions with low data density, score estimates tend to be inaccurate. To address this issue, Vincent (2011) and Song et al. (2021) propose adding noise to the data. When the noise level is sufficiently large, the perturbed data distribution covers the entire space, enabling the training of more robust and scalable score estimators. This principle forms the foundation of our approach. Score-based generative modeling with multi-level noise perturbation allows the construction of reusable motion priors from relatively limited data that lies in a low-dimensional manifold.
@@
Diffusion models approximate a data distribution by progressively corrupting and subsequently denoising samples through a sequence of transformations (Ho et al., 2020). Given a clean sample from the data distribution, the forward process iteratively adds Gaussian noise over $N$ steps to form a Markov chain $\{{\mathbf{x}}^{i}\}_{i=0}^{N}$, defined as:
@@
without performing the full iterative diffusion process. Here, $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$ and $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. This forward diffusion process resembles the multi-level noise perturbation used in score-based generative models (Song et al., 2021). To sample from the learned distribution $p({\mathbf{x}})$, a denoising network $f$ is typically trained using the simple DDPM objective (Ho et al., 2020):
@@
where ${\mathbf{x}}^{i}$ is generated through the forward diffusion process detailed in Equation 3.
Figure 2. Schematic overview of the system. The dashed arrows indicate components used only during policy training. A pretrained motion diffusion model serves as a reusable reward model for motion naturalness via score distillation sampling. The model can be style-conditioned, enabling the policy to learn specific skills or styles without retraining or continuous access to the original motion data.
@@
Once a diffusion model $f$ is trained, samples can be generated by applying the reverse diffusion process or through score distillation sampling (SDS) (Poole et al., 2022). SDS minimizes the KL divergence between the distribution of diffused samples derived from the forward diffusion process $q({\mathbf{x}}^{i}\mid{\mathbf{x}}^{0})$ and the reference data distribution learned by the pretrained diffusion model $f({\mathbf{x}}^{i})$. The gradient of the KL divergence can be estimated using the difference between the score from the forward diffusion process and the prediction from the diffusion model:
@@
where $\mathbf{x}$} is the sample being optimized, and $w(i)$ are coefficients determined by the diffusion noise schedule. Previous work has shown that the weighting function $w(i)$ has limited impact and can be substituted with uniform weighting without negatively affecting performance (Guo et al., 2023). The SDS gradient can also be derived from the diffusion loss in Equation 4, omitting the Jacobian with respect to the score estimator $f$. The SDS loss can be simplified as